OpenStudy (anonymous):

(1+2i)(5-3i) Simplify the given expression. operations on complex numbers please help I tried to do it but I got different answers than the multiple choice @ganeshie8

OpenStudy (anonymous):

the choices are: 2-15i 3+2i 5-6i 11+7i

OpenStudy (anonymous):

@JFraser @jdoe0001

OpenStudy (jfraser):

no clue, sorry. i haven't used complex numbers in forever

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@amistre64 @e.mccormick

OpenStudy (amistre64):

what did you come up with? its just a foil ....

OpenStudy (anonymous):

i got like 3 different answers

OpenStudy (anonymous):

well, i tried it 3 times

OpenStudy (anonymous):

first i got 5+6i

OpenStudy (anonymous):

then idk, then i just did it right now and im at the last step: 5-7i-6i

OpenStudy (amistre64):

1+2i 5-3i ----- 5+10i -3i -6i^2 ----------- 5 +7i -6(-1)

OpenStudy (anonymous):

thats not how u do it

OpenStudy (anonymous):

thats not the FOIL method

OpenStudy (amistre64):

.... that IS the foil method, just not the one you are used to seeing

OpenStudy (anonymous):

with that way, ur answer will be 5+13i , and thats not in the choices

OpenStudy (anonymous):

did u look at the choices above

OpenStudy (amistre64):

5 +7i -6(-1) is not equal to 5+13i

OpenStudy (anonymous):

oh wait woops.. 11 +7i

OpenStudy (anonymous):

i cant do it that way, they didnt show it that way in the lesson so ill nvr do it that way

OpenStudy (anonymous):

 this is what i got when i just did it

OpenStudy (amistre64):

then FOIL it using your way ...

OpenStudy (amistre64):

the trick is in the -6i^2 that results from the Ls

OpenStudy (amistre64):

i^2 = -1 -6i^2 = -6(-1) = 6

OpenStudy (anonymous):

ohhh.. y is it squared

OpenStudy (amistre64):

(1+2i)(5-3i) is what you posted to start with

OpenStudy (anonymous):

yea i know..

OpenStudy (amistre64):

(-3i)(2i) = -6ii = -6i^2

OpenStudy (amistre64):

you forgot to "ii" it

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

-3i * 2i is not equal to 6i

OpenStudy (anonymous):

ok i see now

OpenStudy (amistre64):

err, -6i that is

OpenStudy (anonymous):

can u help with this one

OpenStudy (amistre64):

what do you get when you add the OI parts?

OpenStudy (anonymous):

ok i did it -5+10i+3i-6i^2

OpenStudy (amistre64):

thats good, now 10i + 3i = 13i so the answer will have at least a 13i in it

OpenStudy (anonymous):

then -5+13i-6i^2

OpenStudy (anonymous):

so 1+13i

OpenStudy (anonymous):

can u help with this one